OpenStudy (anonymous):

Ms. Browns class wants to build a rectangular pen for their rabbits. They have 24 ft of fencing . How long should each side be to give the rabbit the most space

OpenStudy (anonymous):

You should draw all of the possible pens and calculate their areas to see which is largest. Example: one pen could be 

OpenStudy (anonymous):

This pen would have a P=24 and A=20...Not the largest possible!

OpenStudy (anonymous):

thats not what the question is asking

OpenStudy (anonymous):

The most space for the rabbit would mean which pen has the greatest area, or square footage.

OpenStudy (anonymous):

ur suppose to write an equation based on the infor but im not sure how

OpenStudy (anonymous):

Are you learning about systems of equations right now?

OpenStudy (anonymous):

ya

OpenStudy (snowfire):

I think this is a related rates problem. You can take the derivative of the Area function and set it equal to 0. This will give you the peak area.

OpenStudy (anonymous):

i thought it could be something like ths 24=12+2x but im unsure

OpenStudy (anonymous):

Solve 24=2l+2w for w... w= (24-2l)/(2) Then substitute this in for the area equation A=l*w A=l*(24-2l)/(2) Graph and look for the l value associated with the max

OpenStudy (anonymous):

Related Rates is the best strategy, (re: Snowfire) if you're in calculus and know how to derive functions.

OpenStudy (anonymous):

i have no clue how to derive functions

OpenStudy (anonymous):

im in precal

OpenStudy (snowfire):

It is as Sam said, you solve for either W or L in terms of the other to make it a one variable problem when you plug it back in. Then you plug that into the area function and solve.

OpenStudy (anonymous):

Well put, thank you Snowfire

OpenStudy (anonymous):

what is the l value

OpenStudy (snowfire):

24=2W+2L 2W=24-2L W=12-L Then you look at the area function: A=L*W Plugging in the new W value from above gives you: A=L*(12-L) A=-L^2+12L Now this is where I would normally use derivatives to finish the problem, but since you aren't expected to know that, you would look at its graph and find the highest point.